OpenStudy (mathmath333):

An orange vendor makes a profit of 20% by selling oranges at a certain price. if he charges Rs. 1.2 higher per orange he would gain 40%. find the original price at which he sold the orange a) 5 b.)4.8 c.) 6 d.)none of these

OpenStudy (mathmath333):

@Michele_Laino

OpenStudy (jhannybean):

Trying to understand how to account for the 40%,

OpenStudy (perl):

Profit = selling price - cost to producer

OpenStudy (perl):

Let x = selling price y = cost to distributor .20x = x - y .40x = (x+1.2) - y

OpenStudy (mathmath333):

y=?

OpenStudy (mathmath333):

u mean \(c\)

OpenStudy (perl):

just to give you an idea of how this comes about, first use a simple example profit = selling price - cost 20 rands = 100 rands - 80 rands

OpenStudy (perl):

oops, rupees

OpenStudy (perl):

20 is 20% of 100

OpenStudy (mathmath333):

by the way option \(b\) and \(c\) both are given wronh

OpenStudy (mathmath333):

*wrong

OpenStudy (perl):

what do you mean

OpenStudy (bibby):

I was feeling D

OpenStudy (mathmath333):

it is giving option \(d\) as right

OpenStudy (bibby):

none of the options increased by 1.2 were a 40% gain anyways

OpenStudy (mathmath333):

there are two selling prices here

OpenStudy (perl):

ok maybe it makes more sense to say profit percent based on cost

OpenStudy (perl):

Let x = original selling price y = cost to producer profit = selling price - cost .2y = x - y .4y = ( x + 1.2) - y

OpenStudy (mathmath333):

\(cool ~thats ~correct\)

OpenStudy (perl):

Check: selling price 7.2 Rs. - cost of 6 Rs. = 1.2 Rs. selling price 8.4 Rs. - cost 6 Rs. = 2.4 Rs. 1.2 Rs. is 20% of 6 Rs. 2.4 Rs. is 40% of 6 Rs.